# Bolzano: Case Studies in LLM-Assisted Mathematical Research

### 18 Apr 2026 v1

@@ Abstract
We report new results on six problems in mathematics and theoretical computer science, produced with the assistance of Bolzano–an open-source multi-agent LLM system. Bolzano orchestrates rounds of interaction between parallel prover agents and a verifier agent while maintaining a persistent knowledge base that is carried across rounds. Classified using the significance–autonomy taxonomy of Feng et al. [F+26b], four of the six results reach the level of publishable research, and three of the six were produced essentially autonomously by Bolzano. Our results provide evidence that LLMs can contribute meaningfully to mathematical research, complementing recent reports by Bubeck et al. [BCE+25], Woodruff et al. [WCAJ+26], and others.
@@ 1 Introduction
This paper provides more evidence in this direction. Bolzano [Bol25b] is an open-source AI system that assists with mathematical research and is available to interested researchers. The system runs research on a problem using several rounds of interaction between several provers and a verifier. Each prover and verifier is implemented by state-of-the-art LLMs (GPT, Gemini, Claude, or others). Its architecture is described in Section 2. We use Bolzano to provide new results for six problems in mathematics and theoretical computer science.
@@
This work has several limitations. First, we do not provide a thorough comparison between Bolzano and single-session chatbot interactions; while such a comparison is feasible in principle, it is difficult to carry out rigorously in a case study with six problems, and the advantage of the multi-agent architecture over direct use of the same underlying models remains unquantified. Second, it is hard to cleanly disentangle the human and AI contributions; even selecting the right problem to work on is a nontrivial human input.
@@
We document six problems in mathematics and theoretical computer science that Bolzano has resolved or made progress on. This was done either autonomously or through collaboration with domain-expert mathematicians and computer scientists. For each of the problems, we include discussion. The corresponding formal results are proved either in the appendices of this paper or in separate papers.
@@
Publishable research Heaps (Section 3.4) PWPP (Section 3.1) KZG (Section 3.3) KKOS Optimization (Section 3.6)
@@
Table 1: Classification of our results by significance and autonomy, following the taxonomy proposed by Feng et al. [F+26b]. to landmark breakthrough. Autonomy ranges from H (primarily human, secondary AI input) through C (human–AI collaboration) to A (essentially autonomous).
@@ 3 Solved Problems
This section contains the six selected problems solved by Bolzano. For each problem, we add discussion about what parts of the research pipeline have been done by a human, and which parts by Bolzano. We always add a link to the Bolzano system containing the proof. However, Bolzano-generated proofs are not intended to be publication-ready. We thus always either add expert-verified proof to appropriate appendix, or link to a paper containing the proof.
@@ 3.1 Black-Box Separation of Adaptive and Non-Adaptive PWPP – Pavel Hubáček
The Bolzano research transcript is available at [Bol26d]. The complete proof appears in [Hub26].
@@ 3.2 Structural Results on Multi-Slope Tilings – Jan Grebík
The proofs appear in Appendix A. The Bolzano research transcript is available at [Bol26e].
@@ 3.3 Special Soundness for Univariate KZG Batching – Pavel Hubáček
The Bolzano research transcript is available at [Bol26b]. The complete proof appears in [BHKM26].
Formal results
Definition 8 (Weak working set property).
Definition 9 (Strong working set property).
@@ 3.5 Partitioning under Function Preimage Constraints – Robert Šámal
The Bolzano research transcript is available at [Bol26c].
@@ 3.6 Complexity of Optimization in KKOS Cultural Dynamics – Martin Koutecký
The proofs appear in Appendix D. The Bolzano research transcript is available at [Bol26a].
@@ Acknowledgements
We thank Tomáš Gavenčiak and Vojtěch Rozhoň for helpful discussions and for their contributions to the development of the Bolzano system. JG, VR, and AZ were supported by the Czech Science Foundation (GA ČR), project No. 26-23599M. RŠ was supported by grant no.25-16627S from the Czech Science Foundation (GAČR). MK was supported by grant no.25-17221S from the Czech Science Foundation (GAČR).

### 24 Apr 2026 v2

@@ Abstract
We report new results on eight problems in mathematics and theoretical computer science, produced with the assistance of Bolzano–an open-source multi-agent LLM system. Bolzano orchestrates rounds of interaction between parallel prover agents and a verifier agent while maintaining a persistent knowledge base that is carried across rounds. Classified using the significance–autonomy taxonomy of Feng et al. [F+26b], six of the eight results reach the level of publishable research, and five of the eight were produced essentially autonomously by Bolzano. Our results provide evidence that LLMs can contribute meaningfully to mathematical research, complementing recent reports by Bubeck et al. [BCE+25], Woodruff et al. [WCAJ+26], and others.
@@ 1 Introduction
This paper provides more evidence in this direction. Bolzano [Bol25b] is an open-source AI system that assists with mathematical research and is available to interested researchers. The system runs research on a problem using several rounds of interaction between several provers and a verifier. Each prover and verifier is implemented by state-of-the-art LLMs (GPT, Gemini, Claude, or others). Its architecture is described in Section 2. We use Bolzano to provide new results for eight problems in mathematics and theoretical computer science.
@@
This work has several limitations. First, we do not provide a thorough comparison between Bolzano and single-session chatbot interactions; while such a comparison is feasible in principle, it is difficult to carry out rigorously in a case study with eight problems, and the advantage of the multi-agent architecture over direct use of the same underlying models remains unquantified. Second, it is hard to cleanly disentangle the human and AI contributions; even selecting the right problem to work on is a nontrivial human input.
@@
We document eight problems in mathematics and theoretical computer science that Bolzano has resolved or made progress on. This was done either autonomously or through collaboration with domain-expert mathematicians and computer scientists. For each of the problems, we include discussion. The corresponding formal results are proved either in the appendices of this paper or in separate papers.
@@
Publishable research Heaps (Section 3.8) PWPP (Section 3.1) KZG (Section 3.3) KKOS Optim. (Section 3.6)
CCE (Section 3.7)
BST Interleaving (Section 3.4)
@@
Table 1: Classification of our results by significance and autonomy, following the taxonomy proposed by Feng et al. [F+26b]. Autonomy ranges from H (primarily human, secondary AI input) through C (human–AI collaboration) to A (essentially autonomous).
@@
    Data structures: We prove that binary search trees have a natural interleaving property, up to $O(\log\log n)$. Bolzano produced the proof autonomously (Section 3.4).
@@
  7. 7.
    Game Theory: We show that, for every $\varepsilon\in(0,1/2)$, there are $n$-player normal-form games where every $\varepsilon$-coarse correlated equilibrium has support of size at least $\Omega\left(\frac{\log{n}}{\varepsilon^{2}\log{(1/\varepsilon)}}\right)$, answering negatively an open problem posed by Babichenko, Barman, and Peretz [BBP14]. This result was proved by Bolzano in four rounds, after a human researcher pointed out the direction of the proof (Section 3.7).
  8. 8.
@@ 3 Solved Problems
This section contains the eight selected problems solved by Bolzano. For each problem, we add discussion about what parts of the research pipeline have been done by a human, and which parts by Bolzano. We always add a link to the Bolzano system containing the proof. However, Bolzano-generated proofs are not intended to be publication-ready. We thus always either add expert-verified proof to appropriate appendix, or link to a paper containing the proof.
@@ 3.1 Black-Box Separation of Adaptive and Non-Adaptive PWPP – Pavel Hubáček
The Bolzano research transcript is available at [Bol26e]. The complete proof appears in [Hub26].
@@ 3.2 Structural Results on Multi-Slope Tilings – Jan Grebík
The proofs appear in Appendix A. The Bolzano research transcript is available at [Bol26f].
@@ 3.3 Special Soundness for Univariate KZG Batching – Pavel Hubáček
The Bolzano research transcript is available at [Bol26c]. The complete proof appears in [BHKM26].
3.4 Interleaving BST Access Sequences and Wilber’s First Bound – Václav Rozhoň
The setting is the classical online binary search tree (BST) problem, in which the cost of an execution equals the total number of nodes touched while serving a sequence of key accesses. The dynamic optimality conjecture of Sleator and Tarjan [ST85] asks whether the splay tree is $O(1)$-competitive with the offline optimum $\mathsf{OPT}(S)$. The best known competitive ratio is $O(\log\log n)$, achieved by Tango trees [DHIP07], whose analysis is driven by Wilber’s first lower bound [Wil89].
A natural question that came up in the work on the dynamic optimality conjecture is what is the behavior of the BST cost under interleaving. For access sequences $X=(x_{1},\dots,x_{m})$ and $Y=(y_{1},\dots,y_{m})$ over $[n]$, write
$X\mathbin{\sqcup\!\sqcup}Y\;:=\;(x_{1},y_{1},x_{2},y_{2},\dots,x_{m},y_{m}).$
A natural conjecture is
$\mathsf{OPT}(X\mathbin{\sqcup\!\sqcup}Y)\;\stackrel{{\scriptstyle?}}{{=}}\;O\bigl(\mathsf{OPT}(X)+\mathsf{OPT}(Y)+m+n\bigr).$
Bolzano proved the exact analogue of this conjecture for Wilber’s first bound.
@@ 3.4 Interleaving BST Access Sequences and Wilber’s First Bound – Václav Rozhoň
Let $Z$ be any two-colored sequence over $[n]$, with monochromatic subsequences $R$ and $B$. Then
$\mathsf{W}(Z)\;\leq\;3\,\mathsf{W}(R)+3\,\mathsf{W}(B)+|Z|.$
In particular, for access sequences $X,Y$ of length $m$,
$\mathsf{W}(X\mathbin{\sqcup\!\sqcup}Y)\;\leq\;3\,\mathsf{W}(X)+3\,\mathsf{W}(Y)+2m.$
Combining Theorem 8 with Wilber’s lower bound and the Tango-tree upper bound $\mathrm{BST\text{-}cost}(S)=O\bigl((\mathsf{W}(S)+|S|)\log\log n+n\bigr)$ yields the following corollary, which says that the BST interleaving conjecture holds up to the standard $O(\log\log n)$ slack separating Wilber’s first bound from the offline optimum.
Corollary 9.
Let $X,Y$ be BST access sequences of length $m$ over $[n]$ that can be served with cost $C_{1}$ and $C_{2}$ respectively. Then $X\mathbin{\sqcup\!\sqcup}Y$ can be served by a BST with cost
$O\bigl((C_{1}+C_{2}+m+n)\log\log n\bigr).$
Both Theorem 8 and Corollary 9 are proved in Appendix F. The Bolzano research transcript is available at [Bol26g].
@@ 3.5 Partitioning under Function Preimage Constraints – Robert Šámal
The Bolzano research transcript is available at [Bol26d].
@@ 3.6 Complexity of Optimization in KKOS Cultural Dynamics – Martin Koutecký
The proofs appear in Appendix D. The Bolzano research transcript is available at [Bol26b].
3.7 Estimating the support size of $\varepsilon$-coarse correlated equilibria – Martin Balko
Nash’s theorem, a fundamental result in game theory, says that every normal-form game $G$ has at least one stable state, called Nash equilibrium. However, all known proofs of this classical result are non-constructive, and there is strong evidence suggesting that there might not be a polynomial-time algorithm to compute Nash equilibria, as this task is PPAD-complete even for 2-player games [CDT09, DGP09]. For this reason, new variants of Nash equilibria that are easier to compute were introduced, such as coarse correlated equilibria.
For a positive integer $n$, let $G=(\{1,\dots,n\},A=A_{1}\times\cdots\times A_{n},u=(u_{1},\dots,u_{n}))$ be a normal-form game of $n$ players with the actions sets $A_{i}$ and utility functions $u_{i}\colon A\to\mathbb{R}$. For $\varepsilon>0$, a probability distribution $P$ on $A$ is an $\varepsilon$-coarse correlated equilibrium if $\mathbb{E}_{a\sim P}[u_{i}(a^{\prime}_{i};a_{-i})-u_{i}(a)]\leq\varepsilon$ for every player $i\in[n]$ and each action $a^{\prime}_{i}\in A_{i}$.
It is significantly simpler to compute Nash equilibria or coarse correlated equilibria with a small support, as those are closer to action profiles, which is witnessed by the classical quasipolynomial algorithm for approximating Nash equilibria by Lipton, Markakis, and Mehta [LMM03]. Thus, there has been a lot of interest in estimating the minimum support size of such equilibria; see, for example [BBP14, Bar18, LMM03]. To capture this formally, we say that a probability distribution $P$ on $A$ is $k$-uniform if there are $k$ action profiles $a^{1},\dots,a^{k}$ such that $P=\frac{1}{k}\sum_{t=1}^{k}a^{t}$.
Babichenko, Barman, and Peretz [BBP14] proved that every normal-form game of $n$ players, each with $m$ actions, admits a $k$-uniform $\varepsilon$-coarse correlated equilibrium for all
$k>\frac{2(\ln{m}+\ln{n})}{\varepsilon^{2}}.$ (1)
They also noted that this bound is asymptotically tight in terms of $m$, but establishing whether this logarithmic in $n$ dependence is tight remained an interesting open problem. Actually, it was not even known whether $k$ cannot be a constant that depends only on $\varepsilon$ and not on $n$. To pinpoint this question, Babichenko, Barman, and Peretz [BBP14] introduced the following problem.
Problem 1 ([BBP14]).
Is there a $k=k(\varepsilon)$ that is independent of $n$, such that in every $n$-player 2-action game there exists $k$-uniform $\varepsilon$-correlated equilibrium?
Balko and Čížek [Bv26] answered this problem negatively by proving the following result with the help of Bolzano.
Theorem 18 ([Bv26]).
There is a constant $C>0$ such that for every $\varepsilon\in(0,1/2)$, there are arbitrarily large integers $n$ and normal-form games $G$ of $n$ players, each with two actions, with payoffs in $\{0,1\}$ such that every $k$-uniform $\varepsilon$-coarse correlated equilibrium in $G$ satisfies
$k\geq\frac{C\cdot\log{n}}{\varepsilon^{2}\log{(1/\varepsilon)}}.$
Note that this result not only shows that $k$ has to depend on $n$, but also provides a lower bound that is asymptotically tight in $n$ as it matches the upper bound (1) in terms of $n$. The full proof appears in [Bv26], and we include it in Appendix E for completeness, as the paper [Bv26] has not yet appeared.
Bolzano worked on this problem for four rounds, guided by a human researcher who pointed out the right research direction. The dependence on $n$ was settled in the first round; the remaining rounds were used to improve the dependence on $\varepsilon$. The Bolzano research transcript is available at [Bol26a].
Formal results
Definition 19 (Weak working set property).
Definition 20 (Strong working set property).
Theorem 21.
@@ Acknowledgements
We thank Tomáš Gavenčiak and Vojtěch Rozhoň for helpful discussions and for their contributions to the development of the Bolzano system. JG, VR, and AZ were supported by the Czech Science Foundation (GA ČR), project No. 26-23599M. RŠ was supported by grant no.25-16627S from the Czech Science Foundation (GAČR). MB and MK were supported by grant no.25-17221S from the Czech Science Foundation (GAČR).